Solve the system of equations in Problem


4x + 3y = 26
3x – 11 y = -7

Answer by anand429(138)


4x + 3y = 26 -(i)
3x – 11 y = -7 -(ii)
Multiplying first eqn by 3,we get,
12x + 9y = 78 -(iii)
Multiplying 2nd eqn by 4,we get,
12x - 44y = -28 -(iv)
Subtracting eqn (iv) from (iii),
53y = 106
or, y =2
putting value of y in eqn (i).
x = 5